OpenStudy (anonymous):

((root1-x^2)/2)-(x(1+x)/2root(1-x^2))=?

OpenStudy (calculusfunctions):

\[\frac{ \sqrt{1-x ^{2}} }{ 2 }-\frac{ x(1+x) }{ 2\sqrt{1-x ^{2}} }=0\]@xkat If this is your equation, then the first step I would take is multiply both sides of the equation by the least common denominator. Thus\[2\sqrt{1-x ^{2}}\left( \frac{ \sqrt{1-x ^{2}} }{ 2 } \right)-2\sqrt{1-x ^{2}}\left( \frac{ x(1+x) }{ 2\sqrt{1-x ^{2}} } \right)=(2\sqrt{1-x ^{2}})(0)\]Now simplify and solve for x. Just remember the restrictions on x.